The management of Priority Health Club claims that its members lose



an average of 10 pounds or more within the first month after joining



the club. A consumer agency that wanted to check this claim took a



random sample of 36 members of this health club and


found that they



lost an average of 9.2 pounds within the first month of membership



with a population standard deviation of 2.4 pounds. Find the


p



value



for this test. What will you decision be if


α



=




.01?




1
Expert's answer
2022-03-07T17:30:09-0500

"H_0: \mu=10."

"H_a:\mu>10."

Test statistic: "z=\frac{9.2-10}{\frac{2.4}{\sqrt{36}}}=-2."

P-value: "p=P(Z>-2)=0.9772."

Since the p-value is greater than 0.01, fail to reject the null hypothesis.